OpenStudy (anonymous):

The equation for the volume of a cone is show below( ill add it in the comments). solve for r

OpenStudy (anonymous):

\[V=\frac{ 1 }{ 3}pir^2h\]

OpenStudy (anonymous):

\[\sqrt{\frac{ 3V }{ \pi*h }}=r\]

OpenStudy (ghazi):

correct